OpenStudy (anonymous):

Use of the Data Booklet is relevant to this question. Anhydrous magnesium nitrate, Mg(NO3)2 , will decompose when heated, giving a white solid and a mixture of two gases X and Y. Y is oxygen. What is the ratio of released of X mass/ released of y released ?

OpenStudy (aaronq):

\(\sf 2 ~Mg(NO_3)_2+\Delta → 2~ MgO + 4 ~NO_2 + ~O_2\) the ratio is based on their stoichiometric coefficients. http://en.wikipedia.org/wiki/Magnesium_nitrate

OpenStudy (anonymous):

why is the answer 1/0.174 ?

OpenStudy (aaronq):

Did you have other values. the ratio should be \(\sf \dfrac{4}{1}\)

OpenStudy (aaronq):

oh wait sorry, it says masses. Use their molar masses

OpenStudy (aaronq):

\(\sf \dfrac{184}{32}=\dfrac{1}{0.17391304}\)

OpenStudy (anonymous):

thank you :)

OpenStudy (aaronq):

no problem !